# A Survey of Zero-Knowledge Proof Based Verifiable Machine Learning

### v1

@@ III-B3 ZKP based Verifiable Inference
Hao et al. [50] conducted an in-depth study on improving the efficiency of ZKP for non-linear functions in ML models. They proposed the use of the table lookup technique and digital decomposition to address the computational cost issues in verifying non-linear functions. Their specific contributions are as follows. (1) Table Lookup Technique: This is one of the core innovations of the paper, aiming to efficiently verify the correctness of non-linear functions through the use of lookup tables. Specifically, complex non-linear functions (such as exponentials, divisions, etc.) are transformed into lookup table problems by storing valid input-output mappings in a lookup table. The prover needs only to prove that the input-output pair exists in the lookup table to complete the function verification. This method avoids the traditional verification approach based on Boolean circuits, eliminating the need to convert arithmetic values into Boolean values, thereby significantly reducing computational complexity. (2) Digital Decomposition: However, directly storing lookup tables of all possible input-output pairs leads to excessively large sizes, especially when the input bit-length is large, causing the size of the lookup table to grow exponentially. To solve the problem of lookup table size explosion, the paper proposes digital decomposition. Large bit-length inputs cause storage requirements and computational costs of lookup tables to increase sharply; therefore, the paper decomposes large bit-length inputs into several smaller numbers (e.g., 5 to 12 bits), effectively reducing the size of the lookup table. For example, an input $x$ can be decomposed as $x=x_{k-1}\parallel\ldots\parallel x_{0}$, where each decomposed number $x_{i}$ has a smaller bit length. This approach allows the lookup table to store mappings of small numbers for verification. Through digital decomposition, the paper supports the verification of large bit-length inputs with less storage overhead and computational cost while maintaining the correctness and completeness of the verification results. Through these two key steps, the authors successfully implemented zero-knowledge proofs for common non-linear functions such as ReLU, Sigmoid, GELU, Softmax, Maxpooling, and normalization. Compared to existing solutions (e.g., Mystique [13]), this research achieved significant performance improvements, with runtime reduced by 50×50$\times$50 × to 179×179$\times$179 ×, and communication costs decreased by 1.2×1.2$\times$1.2 × to 4.8×4.8$\times$4.8 ×. In applicability verification, the framework demonstrated outstanding efficiency advantages in mainstream machine learning tasks (such as inference verification of CNNs and LLMs), especially showing significant performance improvements when handling common functions like ReLU and Softmax. For example, on the ReLU function, runtime was reduced by approximately 100×100$\times$100 ×; on the Softmax function, runtime was reduced by approximately 179×179$\times$179 ×; and on the GELU function, runtime was reduced by approximately $\times7$ to 86×86$\times$86 ×.
Work [48] introduced a ZKP-based verifiable inference framework for deep learning models, aiming to enhance the efficiency and practicality of ZKML while addressing privacy and security concerns in current MLaaS systems. The framework achieves significant improvements in proof efficiency and storage overhead by incorporating sparsity-aware protocols and ternary networks. Its implementation involves the following key steps: (1) Optimized Protocol for Sparse Linear Layers (SpaGKR-LS): The paper proposes an optimized protocol, SpaGKR-LS, specifically designed for the linear layers of neural networks. This protocol achieves proof time that scales linearly with the number of non-zero parameters, significantly reducing computational complexity through techniques such as mode pruning and quantization. By leveraging the sparsity of linear layers, the protocol enables more efficient proofs for sparse networks. (2) Quantization with Ternary Networks: The framework quantizes the model using ternary networks, where parameters are restricted to {-1, 0, 1}. This design eliminates multiplication operations and, by combining sparsity and low-bit-width characteristics, further enhances proof efficiency. The ternary network’s inherent sparsity and simplified arithmetic operations reduce computational overhead, making the proof process more efficient. (3) Modular Framework Design: A modular framework is proposed, compatible with existing ZKML methods based on GKR (Sumcheck Protocol). This modular design allows seamless integration with sparsity-aware protocols, such as the Lasso protocol [71] or the SpaGKR protocol proposed in this work. The flexibility of the modular approach enables efficient verification and adaptability to different use cases. Through these three steps, the proposed framework not only overcomes the efficiency bottlenecks of traditional ZKML methods in handling sparsity and quantized models but also significantly enhances the flexibility and scalability of the framework through its modular design. Experimental results demonstrate the effectiveness of the framework. For sparse linear layers, SpaGKR-LS achieves a 45x speedup in proof time compared to traditional ZKML methods that ignore sparsity. For ternary networks, proof time is further reduced by approximately 5x. Additionally, the framework achieves substantial improvements in verification time and storage efficiency, showcasing its practicality and effectiveness for large-scale deep learning models.
@@
The work of [52] proposes a ZKP-based verifiable inference scheme for deep learning models, aiming to address data leakage and service fraud issues prevalent in current MLaaS platforms. By ensuring the integrity of ML inference processes and results, as well as the privacy and security of ML model parameters, the scheme enhances trust in MLaaS. The proposed scheme integrates non-interactive ZKP with blockchain technology to ensure model integrity verification while protecting sensitive information about model parameters. Its implementation can be divided into three key steps: (1) R1CS Circuit Design for Neural Network Modules: Each module of the neural network—such as convolution, normalization, activation, and pooling layers—is designed as a R1CS circuit to precisely describe and verify its computational logic. To optimize the circuit complexity, several innovations are introduced in the module designs: In the convolutional layers, depthwise separable convolutions are employed instead of traditional convolutions, significantly reducing computational complexity; for the activation functions, minimal polynomial methods are used to approximate complex activation functions (e.g., Swish), avoiding the complexity associated with floating-point operations; in the pooling layers, adaptive average pooling circuits are designed to ensure the circuits can flexibly adapt to different input and output sizes. These optimizations enable the computational logic of each module to be transformed into simple polynomial constraints, significantly reducing the generation time and storage requirements of zk-SNARK proof files. (2) Serial Hash Circuit Connection (SHCC) Algorithm: To address the high complexity introduced by interactions and integrity verification between modules in the modular design of large neural networks, the paper proposes the SHCC algorithm. The large-scale deep learning model is divided into multiple modules, with hash connections ensuring integrity between modules. Specifically, the SHCC algorithm adopts a layer-by-layer hashing method, where each module’s output hash value is generated from the current module’s computation result and the previous module’s hash value, forming a serial hash chain. If the output of any module is tampered with, subsequent module verification will fail due to hash mismatches, thereby ensuring the integrity and security of the entire network. Through these steps, the proposed framework not only ensures the integrity verification of deep learning models but also enhances computational and storage efficiency while guaranteeing privacy protection. Experimental results demonstrate significant optimizations in verification time and storage overhead. Specifically, the verification time is reduced by $54.6\%$. In terms of storage overhead, the transformed R1CS circuits are more concise, reducing the storage requirements of proof files (such as R1CS files, ZKEY files, and WASM files) by $58.1\%$. Notably, the proof files generated using the Groth16 protocol have a fixed size of $4$KB. In the Ethereum environment, verification of proof files is achieved through smart contracts. The verification process is divided into three modules: input module, backbone module, and output module, with proof files generated off-chain separately and verification executed through on-chain smart contracts. Experimental results confirm the feasibility of the scheme through the deployment and execution of smart contracts. Gas consumption tests demonstrate that the verification process is efficient and cost-controllable; for example, the execution cost of the backbone module is 397,434 Gas, and the verification cost is 222,424 Gas.

### v2

@@ III-B3 ZKP-based Verifiable Inference
[44] conducted an in-depth study on improving the efficiency of ZKP for non-linear functions in ML models. They proposed the use of the table lookup technique and digital decomposition to address the computational cost issues in verifying non-linear functions. Their specific contributions are as follows. (1) Table Lookup Technique: This is one of the core innovations of the paper, aiming to efficiently verify the correctness of non-linear functions through the use of lookup tables. Specifically, complex non-linear functions (such as exponentials, divisions, etc.) are transformed into lookup table problems by storing valid input-output mappings in a lookup table. The prover needs only to prove that the input-output pair exists in the lookup table to complete the function verification. This method avoids the traditional verification approach based on Boolean circuits, eliminating the need to convert arithmetic values into Boolean values, thereby significantly reducing computational complexity. (2) Digital Decomposition: However, directly storing lookup tables of all possible input-output pairs leads to excessively large sizes, especially when the input bit-length is large, causing the size of the lookup table to grow exponentially. To solve the problem of lookup table size explosion, the paper proposes digital decomposition. Large bit-length inputs cause storage requirements and computational costs of lookup tables to increase sharply; therefore, the paper decomposes large bit-length inputs into several smaller numbers (e.g., 5 to 12 bits), effectively reducing the size of the lookup table. For example, an input $x$ can be decomposed as $x=x_{k-1}\parallel\ldots\parallel x_{0}$, where each decomposed number $x_{i}$ has a smaller bit length. This approach allows the lookup table to store mappings of small numbers for verification. Through digital decomposition, the paper supports the verification of large bit-length inputs with less storage overhead and computational cost while maintaining the correctness and completeness of the verification results. Through these two key steps, the authors successfully implemented zero-knowledge proofs for common non-linear functions such as ReLU, Sigmoid, GELU, Softmax, Maxpooling, and normalization. Compared to existing solutions (e.g., Mystique  [39]), this research achieved significant performance improvements, with runtime reduced by $50\times$ to $179\times$, and communication costs decreased by $1.2\times$ to $4.8\times$. In applicability verification, the framework demonstrated outstanding efficiency advantages in mainstream machine learning tasks (such as inference verification of CNNs and LLMs), especially showing significant performance improvements when handling common functions like ReLU and Softmax. For example, on the ReLU function, runtime was reduced by approximately $100\times$; on the Softmax function, runtime was reduced by approximately $179\times$; and on the GELU function, runtime was reduced by approximately $77\times$ to $86\times$.
[…]
@@
The work of [54] proposes a ZKP-based verifiable inference scheme for deep learning models, aiming to address data leakage and service fraud issues prevalent in current MLaaS platforms. By ensuring the integrity of ML inference processes and results, as well as the privacy and security of ML model parameters, the scheme enhances trust in MLaaS. The proposed scheme integrates non-interactive ZKP with blockchain technology to ensure model integrity verification while protecting sensitive information about model parameters. Its implementation can be divided into three key steps: (1) R1CS Circuit Design for Neural Network Modules: Each module of the neural network—such as convolution, normalization, activation, and pooling layers—is designed as an R1CS circuit to precisely describe and verify its computational logic. To optimize the circuit complexity, several innovations are introduced in the module designs: In the convolutional layers, depthwise separable convolutions are employed instead of traditional convolutions, significantly reducing computational complexity; for the activation functions, minimal polynomial methods are used to approximate complex activation functions (e.g., Swish), avoiding the complexity associated with floating-point operations; in the pooling layers, adaptive average pooling circuits are designed to ensure the circuits can flexibly adapt to different input and output sizes. These optimizations enable the computational logic of each module to be transformed into simple polynomial constraints, significantly reducing the generation time and storage requirements of zk-SNARK proof files. (2) Serial Hash Circuit Connection (SHCC) Algorithm: To address the high complexity introduced by interactions and integrity verification between modules in the modular design of large neural networks, the paper proposes the SHCC algorithm. The large-scale deep learning model is divided into multiple modules, with hash connections ensuring integrity between modules. Specifically, the SHCC algorithm adopts a layer-by-layer hashing method, where each module’s output hash value is generated from the current module’s computation result and the previous module’s hash value, forming a serial hash chain. If the output of any module is tampered with, subsequent module verification will fail due to hash mismatches, thereby ensuring the integrity and security of the entire network. Through these steps, the proposed framework not only ensures the integrity verification of deep learning models but also enhances computational and storage efficiency while guaranteeing privacy protection. Experimental results demonstrate significant optimizations in verification time and storage overhead. Specifically, the verification time is reduced by $54.6\%$. In terms of storage overhead, the transformed R1CS circuits are more concise, reducing the storage requirements of proof files (such as R1CS files, ZKEY files, and WASM files) by $58.1\%$. Notably, the proof files generated using the Groth16 protocol have a fixed size of 44KB. In the Ethereum environment, verification of proof files is achieved through smart contracts. The verification process is divided into three modules: input module, backbone module, and output module, with proof files generated off-chain separately and verification executed through on-chain smart contracts. Experimental results confirm the feasibility of the scheme through the deployment and execution of smart contracts. Gas consumption tests demonstrate that the verification process is efficient and cost-controllable; for example, the execution cost of the backbone module is 397,434 Gas, and the verification cost is 222,424 Gas.
zkGPT [55] is a recent non-interactive framework for the verifiable inference of large Transformer models such as GPT-2. Building on GKR-style proving, it improves prior systems along three main dimensions: a more efficient matrix-multiplication layer based on grouped multilinear evaluation and precomputation; a “result-as-witness” paradigm for nonlinear and normalization operators such as Softmax, LayerNorm, and GeLU; and system-level optimizations, including constraint fusion and circuit squeezing, that reduce range-check overhead and increase parallelism. The resulting system produces succinct Fiat–Shamir proofs of about 101 KB and reduces end-to-end GPT-2 proving time to under 25 seconds. Although it currently focuses on inference under quantized arithmetic and does not yet address verifiable training or long multi-token generation, zkGPT shows how Transformer-scale inference can be pushed much closer to practical ZKML deployment.